OpenStudy (anonymous):

What is the equation of the line that has a slope of 4/5 and a y-intercept of –1?

OpenStudy (anonymous):

slope intercept again, friend: y=mx+b m=slope b=y-intercept

OpenStudy (anonymous):

i want to help you understand your material! i hope you are not coming here just for answers to your homework!

OpenStudy (anonymous):

its not homework haha and no im not i somewhat understand it better. see im in summer school and tomorrow is my last day and i have to finsih up my final before 1200 am.

OpenStudy (anonymous):

so im kinda rushing through everything i am not very talented at math english and science are more my stronger subjects.

OpenStudy (anonymous):

okay i gotcha man. math and science are my strengths but english, not so much hahah well i hope i could help you understand better

OpenStudy (anonymous):

just memorize slope-intercept form (y=mx+b) and it will answer most of your questions

OpenStudy (anonymous):

You helps a lot thank you so much, and just saying im a girl haha

OpenStudy (anonymous):

oh, well sorry then miss! lol

OpenStudy (anonymous):

what math class are you in? is it high school or?

OpenStudy (anonymous):

haha. now the quesions that look like this how do i answer them

OpenStudy (anonymous):

its high scool and well im going into geometry. Which i find a lot easier than all of this stuff

OpenStudy (anonymous):

okay dont let fractions confuse you! when finding equation of a line given 2 points you will use 2 easy formulas: to find slope: \[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] then you will take what you got for slope and pick one of the 2 original points and use this: \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

so that answer for that equation would be D

OpenStudy (anonymous):

it is actually going to b C

OpenStudy (anonymous):

oh yeah i see what i did wrong.

OpenStudy (anonymous):